# Neural Interpretable Reasoning

### v1

@@ 4 Neural Interpretable Reasoning
  • Semantic transparency: The model must employ high-level, human-understandable concepts, ensuring that the underlying data representations are clear and interpretable.
  • Functional transparency: It is critical to choose a function that maps these concepts to the desired tasks, where both its structure and parameters are comprehensible to the user.
@@
  • Functional mixtures: These help manage the complexity due to a high number of concepts by decomposing the mapping from concepts to tasks into simpler, more interpretable components.
  • Neural re-parametrizations: Both concepts and functions can be neurally re-parametrized allowing to retain the model’s expressivity.
@@
To concretely instantiate our proposal, conside a classification problem where the objective is to predict a target label $Y$ from a set of low-level features (e.g., pixels’ intensities) $X$. Rather than using opaque monolithic model, we propose to leverage expressive power of deep neural networks (DNNs) to generate i) the parameters of a transparent model $W$ and ii) human-understandable data representations $C$ (a.k.a., concepts)—which together form the elements of an interpretable system. The learned transparent model is then symbolically executed to make predictions $Y$:
@@
These two factors represent the neural generation component $\mathbb{P}(C,W\mid X;\theta)$, which re-parametrizes concept representations and functional parameters to ensure expressivity, and the symbolic execution component $\mathbb{P}(Y\mid C;W)$, which guarantees interpretability in the decision-making process.
@@
We refer to the family of models implementing this paradigm as Neural Interpretable Reasoning. This family integrates deep neural network expressivity with the rigorous demands of interpretability by combining semantic transparency, functional transparency, and scalable verification of equivariance. Fig 2 shows a NIR example where a self-driving car must decide whether to brake at an intersection. The architecture first generates both the truth degrees of relevant concepts (e.g., the presence of an ambulance or a green light) and the weights of a simple linear model (e.g., an ambulance is assigned a weight of $2$ because it is positively correlated with braking); then, the linear model is executed on these truth degrees to predict whether to brake. Many well-known XAI techniques can be seen as special cases within this framework. For example, Prototypical Networks (ProtopNets) (Chen et al., 2019), Neural Additive Models (Agarwal et al., 2021), and Concept Bottleneck Models (Koh et al., 2020) all embody aspects of interpretability that align with our proposed approach. More recently, novel approaches such as Concept Memory Reasoning (Debot et al., 2024) and Explanation Bottleneck Models (Yamaguchi & Nishida, 2024) have begun to fully exploit the potential of functional re-parametrization retaining the expressivity of traditional, opaque deep neural networks while supporting the scalable verification of interpretability.

### v2

@@ 4 Neural Interpretable Reasoning
  • Semantic transparency: The model must employ high-level, human-understandable concepts (e.g., as in Kim et al. (2018); Koh et al. (2020); Chen et al. (2020)).
  • Functional transparency: The function that maps these concepts to the desired tasks should have a low-complexity structure (e.g., linear), and its parameters should be interpretable.
@@
  • Functional mixtures: When working in a setup where there is a high number of concepts, functional mixtures help manage the model’s complexity by decomposing the mapping from concepts to tasks into simpler, more interpretable components.
  • Neural re-parametrizations: Both concepts and functions can be neurally re-parametrized, allowing one to retain the model’s expressivity after re-parameterization.
@@
To concretely instantiate our proposal, consider a classification problem where the objective is to predict a target label $Y$ from a set of low-level features (e.g., pixel intensities) $X$. Rather than using an opaque monolithic model, we propose to leverage the expressive power of deep neural networks (DNNs) to generate (i) the parameters of a transparent model $W$, and (ii) human-understandable data representations $C$ (a.k.a., concepts)—which together form the elements of an interpretable system. The learned transparent model is then symbolically executed to make predictions $Y$:
@@
These two factors represent the neural generation component $\mathbb{P}(C,W\mid X;\theta)$, which re-parametrizes concept representations and functional parameters to ensure expressivity, and the symbolic execution component $\mathbb{P}(Y\mid C;W)$, which guarantees interpretability in the decision-making process. We refer to the family of models implementing this paradigm as Neural Interpretable Reasoning. This family integrates deep neural network expressivity with interpretability by combining semantic transparency, functional transparency, and scalable verification of inference equivariance.
@@
Fig 2 shows a NIR example where a self-driving car must decide whether to brake at an intersection. The architecture first generates both the truth degrees of relevant concepts (e.g., the presence of an ambulance or a green light) and the weights of a simple linear model (e.g., an ambulance is assigned a weight of $2$ because it is positively correlated with braking); then, the linear model is executed on these truth degrees to predict whether to brake. Many well-known XAI techniques can be seen as special cases within this framework. For example, Prototypical Networks (ProtopNets) (Chen et al., 2019), Neural Additive Models (Agarwal et al., 2021), and Concept Bottleneck Models (Koh et al., 2020) all embody aspects of interpretability that align with our proposed approach. More recently, novel approaches such as Concept Memory Reasoning (Debot et al., 2024) and Explanation Bottleneck Models (Yamaguchi & Nishida, 2024) have begun to fully exploit the potential of functional re-parametrization retaining the expressivity of traditional, opaque deep neural networks while supporting the scalable verification of interpretability.